# A Survey of World Models for Autonomous Driving

### v3

@@ 4.1 Self-Supervised Learning for World Models
Figure 4: (a) Self-supervised world models lift multi-view images into 3D volumes, forecast future grids, and learn from 2D renders plus occupancy cues, reducing labels (§4.1). (b) Large-scale pre-training on visual point-cloud sequences yields a single backbone that adapts to detection, tracking, mapping, occupancy, and planning (§4.2).
4.2 Large-Scale Pretraining Frameworks
@@
A growing body of research argues that large-scale pre-training of world models is becoming the most economical path to robust autonomous driving. Universal vision/multimodal frameworks [216, 21, 213, 20, 46, 217, 32, 82] ingest millions of image-LiDAR sequences, encode them in a unified BEV or voxel space, and learn to forecast future occupancies or videos. Their strengths lie in rich spatio-temporal context, easy transfer to detection, segmentation, trajectory prediction and even planning, and drastic label-cost reduction. Weaknesses include heavy GPU demand, sensitivity to camera calibration, and lingering reliance on heuristic heads (e.g., simple occupancy decoders). LiDAR/occupancy-centric self-supervision targets pure 3D geometry. Occupancy-MAE [218] masks voxels, AD-L-JEPA [215] reconstructs masked embeddings, and UnO [31] predicts continuous 4D occupancy fields from future point-cloud pseudo-labels. These methods excel at fine-grained geometry, learn on sparse data, and are sensor-agnostic, but they lack rich semantics and require precise temporal alignment. Foundation-level generative world models (e.g., Gaia-1 [15], Gaia-2 [58]) scale to hundreds of millions of frames, synthesize controllable multi-camera scenes, and couple generation with understanding. They promise closed-loop simulation and data augmentation, yet training cost and safety-critical evaluation remain open hurdles.
@@
Collectively, research has progressed from early multimodal prediction networks, through unified 4D pre-training pipelines, to the first attempts at foundation generative models; at each stage the emphasis shifts from “perception only” to “perceive-predict-plan” while annotation demand falls. Hybrid systems are anticipated that fuse dense LiDAR self-supervision with the semantic richness and controllability of vision-centric foundation models. These systems are envisioned to provide task-aware, uncertainty-calibrated ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ software.
4.3 Data Generation for Training
@@
Large-scale autonomous–driving datasets are still dominated by natural logs collected with expensive fleets, yet §3.1 shows that modern world-model generators can manufacture much of what a learner needs. Early image-centric systems [188, 10] could only replay low-resolution clips, but the latest diffusion and Transformer models [15, 58, 46, 14] fuse text, HD-map, trajectory and sensor cues to paint photorealistic, action-conditioned street scenes that retain long-range temporal coherence. Their synthetic videos supplement rare night, rain and construction frames and feed directly into perception or closed-loop policy training. A parallel stream converts multi-view imagery or sparse LiDAR into structured representations before generation. BEV pipelines [33, 31, 22] first lift observations into a bird’s-eye lattice, then roll the lattice forward to harvest scene and intention labels at centimetre accuracy while filtering away distracting appearance noise. Occupancy-grid diffusion [18, 17] and tokenised voxel Transformers [85, 88] push this idea further: they stochastically sculpt 4D grids that obey physics, provide dense semantics and can be queried by planners for risk-aware cost volumes. Finally, LiDAR-native approaches [103, 106] learn score-based or latent denoisers on range images, recreating full sweeps with realistic sensor noise and occlusion artefacts; these meshes or point clouds supply geometry-heavy tasks (e.g., mapping, motion forecasting) without camera bias. Table V summarises representative methods and how their synthetic outputs feed into training.
TABLE V: Usage of data generation4.3).
@@
DriveDreamer-2[37] Generates user-customized videos from text prompts, and the authors discuss adding them to 3D detection training set.
BEVControl[43] BEV sketches are converted to street-view images that fine-tune BEVFormer, boosting NDS by +1.29.
SimGen[47] Simulator + real-layout images serve as augmentation, improving BEV detection & segmentation on nuScenes.
GAIA-2[58] Synthetic multi-camera videos feed Wayve’s offline perception & planning stack for training and safety validation.
@@
GEM[16] Generates scenes from ego trajectory and DINO cues.
Vista[46] High-fidelity video world model evaluated on generation metrics and RL reward.
GAIA-1[15] Demonstrates controllable scene synthesis.
DrivingWorld[64] Video-GPT model measured by FID/FVD.
MiLA[59] Targets long-horizon video synthesis for future training.
CoGen[57] Produces controllable multi-view 3D-consistent scenes.
Figure 5: Application areas and tasks5). (a) Scene understanding fuses multi-camera/LiDAR into a 4D BEV latent that updates every frame, giving precise geometry. (b) Motion prediction rolls latent forward under action cues, outputting collision fields for planners. (c) Simulation turns the generator into an editable 4D world where diffusion seeds rare traffic for safe closed-loop tests. (d) End-to-end driving joins perception, prediction and control in one autoregressive policy.
@@
[…]
    Coverage: By conditioning on text or control commands, generators extrapolate to corner cases absent from logs.
[…]
@@
5 Application Areas and Tasks
@@ 5.1 Scene Understanding
High-fidelity 3D reconstruction is vital, but real sensors are noisy and costly to label. World models answer by compressing multi-camera and LiDAR streams into a unified spatio-temporal latent that continuously refreshes as new frames arrive. Early 4D pre-training shows that one memory state can support detection, mapping and tracking without task-specific heads [21]. Subsequent BEV-centric tokenisers project all modalities into a compact bird’s-eye lattice; decoding that lattice back to pixels or point clouds yields high-resolution semantics and largely removes the need for manual fusion heuristics [32]. A parallel line predicts dense 3D occupancies instead of bounding boxes [90], giving planners free-space priors that remain topologically consistent over seconds rather than frames [18, 85]. Physical realism continues to rise as neural volumetric splats attach radiance and density to each voxel, turning the map into a differentiable oracle for visibility or risk gradients [217, 219]. These advances collapse sensor boundaries, cut annotation budgets, and make scene understanding an anticipatory rather than reactive module; the price is GPU memory and still-uncertain safety certificates for learned geometry.
5.2 Motion Prediction
@@
Future-state inference has moved from extrapolating agent boxes to rolling the latent world forward. Diffusion-based generators evolve the entire scene as a 4D occupancy video that enforces basic physical-consistency losses, allowing planners to clamp candidate ego trajectories and query collision risk in milliseconds [18]. Tokenised video models extend horizons beyond a minute while preserving fine intent cues, widening the temporal window for congestion management and logistics scheduling [220]. Vision-centric hybrids link these 4D forecasts to an end-to-end planner, delivering action-controllable occupancy roll-outs that bypass separate mapping stages [88]. Behaviour realism improves when multi-agent personalities are embedded inside the world; latent “traffic bots” yield socially plausible manoeuvres without brittle lane-change heuristics [198, 197]. The unified latent now integrates map topology, social norms, and kinematics, delivering probability fields that planners can sample or edit instead of juggling disparate trajectory sets. Remaining challenges lie in compute-heavy diffusion roll-outs and the calibration of epistemic uncertainty when imagined futures drift far from training data.
@@
TABLE VI: 4D scene generation comparison6.2). Metrics reported for OccSora [18] vs DynamicCity [17] on CarlaSC [97] and Occ3D-Waymo [83], in both 2D and 3D evaluation spaces.
@@
      CarlaSC       OccSora[18]       16       1.030       28.55       0.008       0.224       0.010       2.257       1559       52.72       0.380       0.151
      DynamicCity[17]       1.040       12.94       0.002       0.307       0.018       2.331       354.2       19.10       0.460       0.170
[…]
      DynamicCity[17]       1.010       36.73       0.001       0.705       0.015       3.206       1806       77.71       0.494       0.026
TABLE VII: Point cloud forecasting performance on OpenScene [221] mini val set (§6.3).
@@
ViDAR [20] 1.34 1.43 1.51 1.60 1.71 1.86 1.58
DFIT-OccWorld-O [86] 0.38 0.72 0.74 0.75 0.79 0.86 0.70
DFIT-OccWorld-V [86] 0.40 0.75 0.78 0.83 0.89 0.90 0.76
5.3 Simulation
Traditional game-engine simulators cannot keep pace with the diversity and rare traffic cases. Generative world models fill that gap by producing controllable, multi-modal replicas of the road. Latent-diffusion frameworks synthesize multi-camera sequences with pixel-level realism and cross-view depth consistency, strong enough for perception-regression tests and data augmentation [15, 46, 58]. Cascade pipelines further blend simulator geometry with real video priors, shrinking the sim-to-real gap and measurably improving BEV detectors when synthetic clips are mixed into training corpora [47, 14]. Geometry-first approaches generate long 4D occupancy volumes, supplying verification tools with dense physical states rather than mere imagery [60, 95, 18]. High-fidelity platforms now close the interaction loop, letting policies influence rendered frames each step and revealing covariate-shift failures before road deployment [29, 206, 203]. However, GPU cost and regulatory acceptance remain hurdles.
5.4 End-to-End Driving
End-to-end pipelines once mapped pixels directly to control signals through black-box regression; world models now provide the memory, foresight and language hooks that make such systems transparent and data-efficient. Generative architectures recast driving as next-token prediction in a multimodal language, producing future frames and waypoint plans that lower collision rates compared with deterministic baselines [22, 127]. Large closed-loop models integrate perception, prediction and control inside a single autoregressive policy, evaluating imagined futures every control cycle to boost rule compliance and ride comfort [196]. Language-vision transformers extend controllability: a single sequence model now obeys natural-language directives like “yield at the zebra crossing, then merge” without extra planners [127, 191]. Lightweight augmentations place a predictive head in front of imitation layers; gradients through imagined roll-outs have been reported to reduce real-world data requirements and expose intermediate occupancy maps for audit [222, 129]. Token limits and long-horizon coherence remain open issues.
6 Performance Comparison
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ auto-
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ dis- ${}_{\!}$ ${}_{\!}$${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ limitations.
6.1 Evaluation Platforms and Benchmarks
CarlaSC. CarlaSC [97] is a synthetic dataset generated via CARLA, offering 4D occupancy at 128×128×$8\!\times\!128\!\times\!$ 128 × 8 resolution within a $25.6m\!\times\!25.6m\!\times 3m$ volume around the ego-vehicle.
NuScenes. nuScenes [36] contains $1,\!000$ scenes split into $700$ train, $150$ val, and $150$ test. It includes six RGB cameras (360) and 32-beam LiDAR with centimeter-level accuracy.
Occ3D-nuScenes. Occ3D-nuScenes [83] is built from nuScenes with $700$ training and $150$ validation sequences, each with $\sim$40 frames at $2$Hz.
Occ3D-Waymo. Occ3D-Waymo [83] is based on Way- mo [38], with $798$ training and $202$ validation sequences, each $\sim$200 frames at $10$Hz, covering $17$ semantic classes.
OpenScene. OpenScene [221], from nuPlan [223], spans over $120$ hours across four cities with $600,\!000{+}$ frames. It uses eight cameras and five LiDARs. OpenScene-mini has $5,\!392$ train and $8,\!729$ val frames.
6.2 4D Scene Generation
Results. To evaluate the effectiveness of DynamicCity [17] in 4D scene generation, comparisons are conducted against OccSora [18] on the Occ3D-Waymo [83] and CarlaSC [97] benchmarks. As reported in Table VI, IS, FID, KID, P, and R are measured in both 2D and 3D spaces. Across both benchmarks and evaluation spaces, DynamicCity consistently delivers higher perceptual quality (IS, FID) and stronger fidelity-diversity trade-offs, validating its effectiveness in generating realistic 4D scenes. Looking forward, research is converging on geometry-consistent 4D diffusion with text/trajectory/physics conditioning and unified RGB-depth-occupancy rendering to support interactive simulators.
TABLE VIII: 4D occupancy forecasting on the Occ3D-nuScenes [83] dataset6.4). Aux. Sup. denotes auxiliary supervision apart from the ego trajectory. Avg. denotes the average performance of that in 1s, 2s, and 3s.

### v4

@@ 4.1. Self-Supervised Learning for World Models
4.2. Large-Scale Pretraining Frameworks
@@
A growing body of research argues that large-scale pre-training of world models is becoming the most economical path to robust autonomous driving. Universal vision/multimodal frameworks (Min et al., 2023, 2024b; Yang et al., 2024e, a; Gao et al., 2024c; Huang et al., 2024a; Zhang et al., 2024a; Li et al., 2025b) ingest millions of image-LiDAR sequences, encode them in a unified BEV or voxel space, and learn to forecast future occupancies or videos. Their strengths lie in rich spatio-temporal context, easy transfer to detection, segmentation, trajectory prediction and even planning, and drastic label-cost reduction. Weaknesses include heavy GPU demand, sensitivity to camera calibration, and lingering reliance on heuristic heads (e.g., simple occupancy decoders). LiDAR/occupancy-centric self-supervision targets pure 3D geometry. Occupancy-MAE (Min et al., 2024a) masks voxels, AD-L-JEPA (Zhu et al., 2025a) reconstructs masked embeddings, and UnO (Agro et al., 2024) predicts continuous 4D occupancy fields from future point-cloud pseudo-labels. These methods excel at fine-grained geometry, learn on sparse data, and are sensor-agnostic, but they lack rich semantics and require precise temporal alignment. Foundation-level generative world models (e.g., Gaia-1 (Hu et al., 2023a), Gaia-2 (Russell et al., 2025)) scale to hundreds of millions of frames, synthesize controllable multi-camera scenes, and couple generation with understanding. They promise closed-loop simulation and data augmentation, yet training cost and safety-critical evaluation remain open hurdles.
@@
Research has progressed from early multimodal prediction networks, through unified 4D pre-training pipelines, to the first attempts at foundation generative models; at each stage the emphasis shifts from “perception only” to “perceive-predict-plan” while annotation demand falls. Hybrid systems are anticipated that fuse dense LiDAR self-supervision with the semantic richness and controllability of vision-centric foundation models. These systems are envisioned to provide task-aware, uncertainty-calibrated priors that can be fine-tuned online. Several open challenges remain: scalable temporal memory, principled safety validation, and energy-efficient training. If these hurdles are overcome, world-model pre-training is likely to become the standard backbone for all levels of autonomous-driving software.
4.3. Data Generation for Training
@@
Large-scale autonomous–driving datasets are still dominated by natural logs collected with expensive fleets, yet §3.1 shows that modern world-model generators can manufacture much of what a learner needs. Early image-centric systems (Kim et al., 2021; Wang et al., 2024c) could only replay low-resolution clips, but the latest diffusion and Transformer models (Hu et al., 2023a; Russell et al., 2025; Gao et al., 2024c; Li et al., 2024d) fuse text, HD-map, trajectory and sensor cues to paint photorealistic, action-conditioned street scenes that retain long-range temporal coherence. Their synthetic videos supplement rare night, rain and construction frames and feed directly into perception or closed-loop policy training. A parallel stream converts multi-view imagery or sparse LiDAR into structured representations before generation. BEV pipelines (Hu et al., 2021b; Agro et al., 2024; Zheng et al., 2025b) first lift observations into a bird’s-eye lattice, then roll the lattice forward to harvest scene and intention labels at centimetre accuracy while filtering away distracting appearance noise. Occupancy-grid diffusion (Wang et al., 2024b; Bian et al., 2025) and tokenised voxel Transformers (Zheng et al., 2025a; Yang et al., 2025c) push this idea further: they stochastically sculpt 4D grids that obey physics, provide dense semantics and can be queried by planners for risk-aware cost volumes. Finally, LiDAR-native approaches (Zyrianov et al., 2022; Hu et al., 2024b) learn score-based or latent denoisers on range images, recreating full sweeps with realistic sensor noise and occlusion artefacts; these meshes or point clouds supply geometry-heavy tasks (e.g., mapping, motion forecasting) without camera bias. Table 5 summarises representative methods and how their synthetic outputs feed into training.
Table 5. Usage of data generation4.3).
@@
    DriveDreamer-2(Zhao et al., 2025b)      Generates user-customized videos from text prompts, and the authors discuss adding them to 3D detection training set.
    BEVControl(Yang et al., 2023b)      BEV sketches are converted to street-view images that fine-tune BEVFormer, boosting NDS by +1.29.
    SimGen(Zhou et al., 2024)      Simulator + real-layout images serve as augmentation, improving BEV detection & segmentation on nuScenes.
    GAIA-2(Russell et al., 2025)      Synthetic multi-camera videos feed Wayve’s offline perception & planning stack for training and safety validation.
@@
    GEM(Hassan et al., 2024)      Generates scenes from ego trajectory and DINO cues.
    Vista(Gao et al., 2024c)      High-fidelity video world model evaluated on generation metrics and RL reward.
    GAIA-1(Hu et al., 2023a)      Demonstrates controllable scene synthesis.
    DrivingWorld(Hu et al., 2024a)      Video-GPT model measured by FID/FVD.
    MiLA(Wang et al., 2025a)      Targets long-horizon video synthesis for future training.
    CoGen(Ji et al., 2025)      Produces controllable multi-view 3D-consistent scenes.
@@
  • Coverage: By conditioning on control commands, generators extrapolate to corner cases absent from logs.
@@
5. Application Areas and Tasks
@@ 5.1. Scene Understanding
High-fidelity 3D reconstruction is vital, but real sensors are noisy and costly to label. World models answer by compressing multi-camera and LiDAR streams into a unified spatio-temporal latent that continuously refreshes as new frames arrive. Early 4D pre-training shows that one memory state can support detection, mapping and tracking without task-specific heads (Min et al., 2024b). Subsequent BEV-centric tokenisers project all modalities into a compact bird’s-eye lattice; decoding that lattice back to pixels or point clouds yields high-resolution semantics and largely removes the need for manual fusion heuristics (Zhang et al., 2024a). A parallel line predicts dense 3D occupancies instead of bounding boxes (Zuo et al., 2025), giving planners free-space priors that remain topologically consistent over seconds rather than frames (Wang et al., 2024b; Zheng et al., 2025a). Physical realism continues to rise as neural volumetric splats attach radiance and density to each voxel, turning the map into a differentiable oracle for visibility or risk gradients (Huang et al., 2024a; Kerbl et al., 2023). These advances collapse sensor boundaries, cut annotation budgets, and make scene understanding an anticipatory rather than reactive module; the price is GPU memory and still-uncertain safety certificates for learned geometry.
Figure 5. Application areas and tasks5). (a) Scene understanding fuses multi-camera/LiDAR into a 4D BEV latent that updates every frame, giving precise geometry. (b) Motion prediction rolls latent forward under action cues, outputting collision fields for planners. (c) Simulation turns the generator into an editable 4D world where diffusion seeds rare traffic for safe closed-loop tests. (d) End-to-end driving joins perception, prediction and control in one autoregressive policy.
5.2. Motion Prediction
@@
Future-state inference has moved from extrapolating agent boxes to rolling the latent world forward. Diffusion-based generators evolve the entire scene as a 4D occupancy video that enforces basic physical-consistency losses, allowing planners to clamp candidate ego trajectories and query collision risk in milliseconds (Wang et al., 2024b). Tokenised video models extend horizons beyond a minute while preserving fine intent cues, widening the temporal window for congestion management and logistics scheduling (Guo et al., 2024). Vision-centric hybrids link these 4D forecasts to an end-to-end planner, delivering action-controllable occupancy roll-outs that bypass separate mapping stages (Yang et al., 2025c). Behaviour realism improves when multi-agent personalities are embedded inside the world; latent “traffic bots” yield socially plausible manoeuvres without brittle lane-change heuristics (Zhang et al., 2023a; Suo et al., 2021). The unified latent now integrates map topology, social norms, and kinematics, delivering probability fields that planners can sample or edit instead of juggling disparate trajectory sets. Remaining challenges lie in compute-heavy diffusion roll-outs and the calibration of epistemic uncertainty when imagined futures drift far from training data.
5.3. Simulation
Traditional game-engine simulators cannot keep pace with the diversity and rare traffic cases. Generative world models fill that gap by producing controllable, multi-modal replicas of the road. Latent-diffusion frameworks synthesize multi-camera sequences with pixel-level realism and cross-view depth consistency, strong enough for perception-regression tests and data augmentation (Hu et al., 2023a; Gao et al., 2024c; Russell et al., 2025). Cascade pipelines further blend simulator geometry with real video priors, shrinking the sim-to-real gap and measurably improving BEV detectors when synthetic clips are mixed into training corpora (Zhou et al., 2024; Li et al., 2024d). Geometry-first approaches generate long 4D occupancy volumes, supplying verification tools with dense physical states rather than mere imagery (Guo et al., 2025; Gu et al., 2024; Wang et al., 2024b). High-fidelity platforms now close the interaction loop, letting policies influence rendered frames each step and revealing covariate-shift failures before road deployment (Yang et al., 2024d; Yan et al., 2025b; Yang et al., 2023a). However, GPU cost and regulatory acceptance remain hurdles.
5.4. End-to-End Driving
End-to-end pipelines once mapped pixels directly to control signals through black-box regression; world models now provide the memory, foresight and language hooks that make such systems transparent and data-efficient. Generative architectures recast driving as next-token prediction in a multimodal language, producing future frames and waypoint plans that lower collision rates compared with deterministic baselines (Zheng et al., 2025b; Chen et al., 2024). Large closed-loop models integrate perception, prediction and control inside a single autoregressive policy, evaluating imagined futures every control cycle to boost rule compliance and ride comfort (Zheng et al., 2024). Language-vision transformers extend controllability: a single sequence model now obeys natural-language directives like “yield at the zebra crossing, then merge” without extra planners (Chen et al., 2024; Tan et al., 2023). Lightweight augmentations place a predictive head in front of imitation layers; gradients through imagined roll-outs have been reported to reduce real-world data requirements and expose intermediate occupancy maps for audit (Li et al., 2025a, d). Token limits and long-horizon coherence remain open issues.
6. Performance Comparison
This section presents an evaluation of world models for autonomous driving across tasks and metrics. Based on discussions in §5, representative algorithms are benchmarked to provide evidence of their strengths and limitations.
6.1. Evaluation Platforms and Benchmarks
CarlaSC. CarlaSC (Wilson et al., 2022) is a synthetic dataset generated via CARLA, offering 4D occupancy at $128\!\times\!128\!\times\!8$ resolution within a $25.6m\!\times\!25.6m\!\times 3m$ volume around the ego-vehicle.
NuScenes. nuScenes (Caesar et al., 2020) contains 1,0001,\!000 scenes split into $700$ train, $150$ val, and $150$ test. It includes six RGB cameras (360) and 32-beam LiDAR with centimeter-level accuracy.
Occ3D-nuScenes. Occ3D-nuScenes (Tian et al., 2024) is built from nuScenes with $700$ training and $150$ validation sequences, each with $\sim$40 frames at 22Hz.
Occ3D-Waymo. Occ3D-Waymo (Tian et al., 2024) is based on Waymo (Sun et al., 2020), with $798$ training and $202$ validation sequences, each $\sim$200 frames at 1010Hz, covering $17$ semantic classes.
OpenScene. OpenScene (Zhang et al., 2024e), from nuPlan (Cao and De Charette, 2022), spans over $120$ hours across four cities with 600,000+600,\!000{+} frames. It uses eight cameras and five LiDARs. OpenScene-mini has 5,3925,\!392 train and 8,7298,\!729 val frames.
6.2. 4D Scene Generation
Results. To evaluate the effectiveness of DynamicCity (Bian et al., 2025) in 4D scene generation, comparisons are conducted against OccSora (Wang et al., 2024b) on the Occ3D-Waymo (Tian et al., 2024) and CarlaSC (Wilson et al., 2022) benchmarks. As reported in Table 6, IS, FID, KID, P, and R are measured in both 2D and 3D spaces. Across both benchmarks and evaluation spaces, DynamicCity consistently delivers higher perceptual quality (IS, FID) and stronger fidelity-diversity trade-offs, validating its effectiveness in generating realistic 4D scenes. Looking forward, research is converging on geometry-consistent 4D diffusion with text/trajectory/physics conditioning and unified RGB-depth-occupancy rendering to support interactive simulators.
@@ 6.1. Evaluation Platforms and Benchmarks
Table 6. 4D scene generation comparison6.2). Metrics reported for OccSora (Wang et al., 2024b) vs DynamicCity (Bian et al., 2025) on CarlaSC (Wilson et al., 2022) and Occ3D-Waymo (Tian et al., 2024), in both 2D and 3D evaluation spaces.
@@
CarlaSC OccSora(Wang et al., 2024b) 16 1.030 28.55 0.008 0.224 0.010 2.257 1559 52.72 0.380 0.151
DynamicCity(Bian et al., 2025) 1.040 12.94 0.002 0.307 0.018 2.331 354.2 19.10 0.460 0.170
[1pt/1pt] Occ3D-Waymo OccSora(Wang et al., 2024b) 16 1.005 42.53 0.049 0.654 0.004 3.129 3140 12.20 0.384 0.001
DynamicCity(Bian et al., 2025) 1.010 36.73 0.001 0.705 0.015 3.206 1806 77.71 0.494 0.026
Table 7. Point cloud forecasting performance on OpenScene (Zhang et al., 2024e) mini val set (§6.3).
@@
    ViDAR (Yang et al., 2024a)     1.34     1.43     1.51     1.60     1.71     1.86     1.58
    DFIT-OccWorld-O (Zhang et al., 2024c)     0.38     0.72     0.74     0.75     0.79     0.86     0.70
    DFIT-OccWorld-V (Zhang et al., 2024c)     0.40     0.75     0.78     0.83     0.89     0.90     0.76
6.3. Point Cloud Forecasting
Results. Table 7 shows that both DFIT‐OccWorld variants substantially outperform the ViDAR baseline at every forecast horizon. DFIT-OccWorld-O trims mean CD to 0.70m2 and stays ahead of ViDAR across all horizons; the V-variant follows at 0.76m2, and the gap widens as lead time grows. The field is shifting from range-image CNN/LSTM to token-based diffusion, leveraging image priors and physics constraints to sustain long-horizon precision.
Table 8. 4D occupancy forecasting on the Occ3D-nuScenes (Tian et al., 2024) dataset6.4). Aux. Sup. denotes auxiliary supervision apart from the ego trajectory. Avg. denotes the average performance of that in 1s, 2s, and 3s.